How do I do Elimination using multiplication. I'm in the 9th grade my book is the Glencoe Mathematics Algebra 1. Can you please help me and e-mailme back explaining this to me in detail thanks ALOT this website is very helpful and agai thanks alot. from Christina 1-18-06 PLEASE EXPLAIN VERY DETAILED thanks.

Answer by elima(1433)

To solve asystem of equations you can addor subtract the two equations. You do this to "Eliminate" one variable and solvethe equation for the other. When you have solved the equation for the varialbe, then you can plug the answer into one of the priginal equation and solve for the other variable.
ex.
Lets say we are given these two equations and are told to find x& y;
2x+2y=48
x-2y=12
To do this we begin by adding the two equations together, and when you do this the y's cancel each other out;
2x+2y=48
x-2y=12
---------
3x =60
Now you can solve for x,
3x=60
x=20
Now that youhave the answer for x, you can plug that answer into the priginal equation and solve for y;
2(20)+2y=48
40+2y=48
2y=48-40
2y=8
y=4
I hope you are able to understand this explanation.
=)